# Efficient Parameter Calibration of Numerical Weather Prediction Models via Evolutionary Sequential Transfer Optimization

### v1

@@ III-C Representation Extractor
where $N$ represents the number of samples in a training batch, and $\|\cdot\|_{2}$ denotes the $L_{2}$ norm. By minimizing the loss function $\mathcal{L}$, the model compels the encoder to retain crucial feature information from the original data within the low-dimensional latent vector $z_{t}$, thereby achieving high-quality reconstruction.
@@ IV-A Experimental Settings
    where $s_{i,t}$ and $\hat{s}_{i,t}$ represent the simulated value and the observed value, respectively, for the $i$-th grid cell at time step $t$, and $N$ denotes the total number of grid cells.
@@ IV-C Performance Comparison of Calibration Algorithms
Specific settings were adopted for this comparative study. During the initialization phase, both KMO and MOASMO require Quasi-Monte Carlo (QMC) sampling to generate parameter vectors ($20$ samples, identical for both algorithms), which are subsequently evaluated by WRF to construct the surrogate models for the problem. These methods solve each calibration task independently, without transferring knowledge from the source task archive. Furthermore, by calculating the similarity between the $10$ target tasks and the $20$ source tasks, we observed that Target Tasks 1-7 could identify highly similar tasks within the source archive, whereas Target Tasks 8-10 could not. This distribution aligns with realistic calibration scenarios.
The comparative results are presented in Table III. Here, $\Delta HV(\%)$ denotes the percentage difference in HV values between the baselines (KMO and MOASMO) and SEETO at the $20$-th evaluation; $Add.FE(\%)$ represents the additional percentage of total evaluation steps required for KMO and MOASMO to achieve the same HV value that SEETO attained at the $20$-th step (relative to the base $20$ steps). In scenarios characterized by high similarity between the target and source tasks (Tasks 1-7), SEETO demonstrated a distinct computational advantage under the condition of $20$ WRF evaluations. Regardless of the varying difficulty across tasks, SEETO consistently maintained the lead. Taking Task 4 as an example, the performance gap between the baselines and SEETO reached as high as $28.88\%$. This indicates that within complex parameter calibration landscapes, relying solely on QMC sampling makes it difficult to capture high-quality solutions. In contrast, by integrating historical populations and surrogate models into the optimization process of the current target task, SEETO successfully located superior parameter regions via transferred knowledge. Crucially, the additional computational cost required to bridge these HV gaps is prohibitive. For Task 4, MOASMO required at least an additional $200\%$ of evaluation steps to match the performance level of SEETO at step $20$. Even for Task 2, where the gap was smaller, MOASMO still necessitated an expenditure of $105\%$ more computational resources.

### v2

@@ III-C Representation Extractor
where NN represents the number of samples in a training batch, and $\|\cdot\|_{2}$ denotes the $L_{2}$ norm. By minimizing the loss function $\mathcal{L}$, the model compels the encoder to retain crucial feature information from the original data within the low-dimensional latent vector $z_{t}$, thereby achieving high-quality reconstruction.
@@ IV-A Experimental Settings
    where $s_{i,t}$ and $\hat{s}_{i,t}$ represent the simulated value and the observed value, respectively, for the $i$-th grid cell at time step $t$, and NN denotes the total number of grid cells.
@@ IV-C Performance Comparison of Calibration Algorithms
Specific settings were adopted for this comparative study. During the initialization phase, both KMO and MOASMO require Quasi-Monte Carlo (QMC) sampling to generate parameter vectors (2020 samples, identical for both algorithms), which are subsequently evaluated by WRF to construct the surrogate models for the problem. These methods solve each calibration task independently, without transferring knowledge from the source task archive. Furthermore, by calculating the similarity between the 1010 target tasks and the 2020 source tasks, we observed that Target Tasks 1-7 could identify highly similar tasks within the source archive, whereas Target Tasks 8-10 could not. This distribution aligns with realistic calibration scenarios.
The comparative results are presented in Table III. Here, $\Delta HV(\%)$ denotes the percentage difference in HV values between the baselines (KMO and MOASMO) and SEETO at the 2020-th evaluation; $Add.FE(\%)$ represents the additional percentage of total evaluation steps required for KMO and MOASMO to achieve the same HV value that SEETO attained at the 2020-th step (relative to the base 2020 steps). In scenarios characterized by high similarity between the target and source tasks (Tasks 1-7), SEETO demonstrated a distinct computational advantage under the condition of 2020 WRF evaluations. Regardless of the varying difficulty across tasks, SEETO consistently maintained the lead. Taking Task 4 as an example, the performance gap between the baselines and SEETO reached as high as $28.88\%$. This indicates that within complex parameter calibration landscapes, relying solely on QMC sampling makes it difficult to capture high-quality solutions. In contrast, by integrating historical populations and surrogate models into the optimization process of the current target task, SEETO successfully located superior parameter regions via transferred knowledge. Crucially, the additional computational cost required to bridge these HV gaps is prohibitive. For Task 4, MOASMO required at least an additional $200\%$ of evaluation steps to match the performance level of SEETO at step 2020. Even for Task 2, where the gap was smaller, MOASMO still necessitated an expenditure of $105\%$ more computational resources.